The measure of an angle is 28 degrees more than its complement. Find the measure of each angle.
Answer by mananth(16949)

let angle be x
complement = 90-x
x=(90-x)+28
x=90-x+28
add x to both sides
x+x=90-28
2x=62
x= 62/2
x=31 deg